# AI for Science

### v1

@@ 5 Conclusion & Future Work
Future work will focus on integrating experimental feedback into the reward loop, incorporating structural modeling and docking-based scores, and expanding reasoning depth to capture multi-position or scaffold-level edits[46, 12]. Addressing these directions will move PepThink-R1 closer to a practical assistant for therapeutic peptide discovery, bridging symbolic reasoning with data-driven optimization.
@@ B.1 Implementation Details
For the training of PepThink-R1, we used a chemical knowledge-based LLM LlaSMol [31] as the base model which was derived from Mistral-7b-v0.1. We designed the prompt template to clearly specify the required output structure, which included using <think> and <SMILES> tags to separate the reasoning process and the final output SMILES. The detailed prompt is shown in Table. 1. For SFT training hyperparameters, we adopted a cutoff length of 2048 tokens, trained for 3 epochs with a learning rate of $5\times 10^{-5}$, batch size of 4, and gradient accumulation of 8 steps. We used cosine learning rate scheduling with 20 warmup steps, and set the maximum number of training samples to 26,000 with 10% held out for validation. Checkpoints were saved every 50 steps with logging performed every 5 steps. LoRA [47] was applied with rank 8, scaling factor 16, and dropout 0.05. We adopted LLama-Factory [48] as the SFT training framework. For RL training hyperparameters, we set training batch size to be 128, with 8 rollouts generated for each sample. The sampling temperature was set to 1.0 to encourage response diversity, and optimization was done with an initial learning rate of $1\times 10^{-6}$. The KL divergence coefficient of GRPO loss was fixed to be $1\times 10^{-3}$ throughout the training process. We adopted verl [49] as the RL training framework. All training was conducted on 8 NVIDIA Tesla H100 80GB GPUs.

### v2

@@ 5 Conclusion & Future Work
Future work will focus on integrating experimental feedback into the reward loop, incorporating structural modeling and docking-based scores, and expanding reasoning depth to capture multi-position or scaffold-level edits[liu2025acereason, guo2025deepseek]. Addressing these directions will move PepThink-R1 closer to a practical assistant for therapeutic peptide discovery, bridging symbolic reasoning with data-driven optimization.
6 Acknowledgment
The authors thank our colleagues Ruibo Zhang, Yixiang Mao, Marissa Vavrek, Tanmoy Pal, Joseph Forbes for helpful discussions on peptide design, property prediction, and troubleshooting.
7 Declaration of Competing Interest
All authors, when performing the research, worked for Merck Sharp & Dohme LLC, a subsidiary of Merck & Co., Inc., Rahway, NJ, USA. Specifically, R.W., H.Z., and T.N. were summer interns, and the others worked as full-time employees.
8 Data and Code Availability
Raw data is downloaded from CycPeptMPDB database and available at http://cycpeptmpdb.com/. CycloPs is downloaded from https://github.com/fergaljd/cyclops. LLaMA-Factory is used for SFT and downloaded from https://github.com/hiyouga/LLaMA-Factory/tree/v0.9.2. verl is used for RL and downloaded from https://github.com/volcengine/verl. PepThink-R1 code is available at https://github.com/MSDLLCpapers/PepThink-R1.
@@ B.1 Implementation Details
For the training of PepThink-R1, we used a chemical knowledge-based LLM LlaSMol [yu2024llasmol] as the base model which was derived from Mistral-7b-v0.1. We designed the prompt template to clearly specify the required output structure, which included using <think> and <SMILES> tags to separate the reasoning process and the final output SMILES. The detailed prompt is shown in Table. 1. For SFT training hyperparameters, we adopted a cutoff length of 2048 tokens, trained for 3 epochs with a learning rate of $5\times 10^{-5}$, batch size of 4, and gradient accumulation of 8 steps. We used cosine learning rate scheduling with 20 warmup steps, and set the maximum number of training samples to 26,000 with 10% held out for validation. Checkpoints were saved every 50 steps with logging performed every 5 steps. LoRA [hu2022lora] was applied with rank 8, scaling factor 16, and dropout 0.05. We adopted LLama-Factory [zheng2024llamafactory] as the SFT training framework. For RL training hyperparameters, we set training batch size to be 128, with 8 rollouts generated for each sample. The sampling temperature was set to 1.0 to encourage response diversity, and optimization was done with an initial learning rate of $1\times 10^{-6}$. The KL divergence coefficient of GRPO loss was fixed to be $1\times 10^{-3}$ throughout the training process. We adopted verl [sheng2024hybridflow] as the RL training framework. All training was conducted on 8 NVIDIA Tesla H100 80GB GPUs.